OpenStudy (anonymous):

ok i need help with this one i have trouble simplifying. simplify the following expression: 121 divided by 11+3*4 divided by 2. this is all one expression but i dont know what use as the dived sign. A. 15 B.28 C.17 D. 4.32 @jim_thompson5910 @Nnesha

OpenStudy (anonymous):

i got 17

OpenStudy (anonymous):

17 is the correct answer

jimthompson5910 (jim_thompson5910):

I agree: 121/11+3*4/2 = 17 you use a forward slash to mean divide

OpenStudy (anonymous):

ok because i dont want the / to make people think its a fraction or something.

jimthompson5910 (jim_thompson5910):

division and fractions are one in the same really eg: one-half = 1/2 = 0.5

OpenStudy (adi3):

I got 2.6

OpenStudy (adi3):

is that right